OpenStudy (anonymous):

Solve for x in the first equation, then solve the system using substitution. Write your answer as (x,y) 2x-6y=14 4x-y=9

OpenStudy (whpalmer4):

\[2x-6y=14\]\[4x-y=9\] Okay, solve the first equation for \(x\): \[2x-6y=14\]add \(6y\) to both sides\[2x-6y+6y=14+6y\]\[2x=14+6y\]Now divide both sides by \(2\): \[\frac{2x}{2} = \frac{14+6y}{2}\]\[x = 7+3y\] Now replace \(x\) in \(4x-y=9\) with \((7+3y)\) and solve for the value of \(y\). When you have the value of \(y\), plug it into \(x = 7+3y\) and find \(x\). Report your answer as instructed.